# Multispecies inhomogeneous $t$-PushTASEP from antisymmetric fusion

### v2

@@ 5.3. Wanted terms
This subsection and the next form the technical focus of the proof. From the definition of $w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)$ around (11), the wanted terms generally correspond to the situation where the minimum $h_{0}$ of the moved particle types $(h_{0},\ldots,h_{g})$ in Lemma 5 is zero, i.e., $h_{0}=0$. Then (63) is written down explicitly as
$\prod_{i=1}^{g}\frac{(1-t)t^{\ell_{h_{i}}}}{1-t^{K_{h_{i}}}}=D_{\bf m}^{-1}\sum_{k=d}^{n+1}(-1)^{k-1}x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}.$ (64)
Our calculation of the RHS of (64) consists of two steps.
@@
Step 1. We consider the “leading term” $k=d$ in the RHS of (63) and the corresponding reduced diagram, in which the carriers are uniquely determined, as shown in Example 6 (b) and (c). We claim that
$x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{d}(0)|\boldsymbol{\sigma}\rangle_{j}=(-1)^{d-1}(1-t)\prod_{i=1}^{g}(1-t)t^{\ell_{h_{i}}},$ (65)
@@
where $\ell_{h}$ has been defined prior to (11). Let us justify the origin of the constituent factors (i) sign, (ii) powers of $(1-t)$, (iii) powers of $t$, individually.
@@
(i) The sign of a vertex can become negative only for non-diagonal transitions, which occur at the $g+1$ vertices in the reduced diagram (59). From the comment following (33) and the conditions in Lemma 5, the $g+1$ vertices corresponding to the vertical arrows $r_{i}\rightarrow s_{i}$ in (60b) have $+$ signs for $i=1,\ldots,g$ and $(-1)^{d-1}$ for $i=0$.
(ii) From Table 1, the contributions of $(1-t)$ at each of the $g+1$ vertices results in a factor of $(1-t)^{g+1}$.
@@
(iii) From Table 1, the power of $t$ can be evaluated as the sum of the quantities of the form $a_{c+1}+\cdots+a_{n}$ in the multiplicity representation of the carriers ${\bf a}=(a_{0},\ldots,a_{n})$, attached to each vertex. This formula implies that a particle of type $h$ in the carriers contributes $[c<h]\in\{0,1\}$ whenever it passes over a site $i$ occupied with $\sigma_{i}=c$. Alternatively, this can be calculated as the total contribution collected by the moved particles $h_{0},\ldots,h_{g}$ from the smaller-species particles in $\boldsymbol{\sigma}$. This precisely leads to $\ell_{h_{1}}+\cdots+\ell_{h_{g}}$, where $\ell_{h_{0}}$ can be excluded due to $\ell_{h_{0}}=\ell_{0}=0$. Thus the factor $\prod_{i=1}^{g}t^{\ell_{h_{i}}}$ is obtained as claimed. The reformulation in the calculation described here is analogous to the transition from the Eulerian picture, which tracks properties at fixed spatial points, to the Lagrangian picture, which follows individual particles, in fluid mechanics. In our context, it also incorporates the contribution from the vertices corresponding to the diagonal transitions efficiently via the quantities $\ell_{h}$’s.
@@
Step 2. Let us turn to the $k=d+1,\ldots,n+1$ terms in (63). We illustrate the idea of evaluating them along Example 6 (b) for $k=5$ and $n=7$ $(d=2,g=3)$. The carriers from $\mathscr{B}^{5}$ are no longer unique. However, those satisfying the weight conservation with $\boldsymbol{\sigma}$ and $\boldsymbol{\sigma}^{\prime}$ are exactly those obtained just by supplementing the common three letters from the yet unused ones $\{1,5,6,7\}$ to the existing ones everywhere. For instance, choosing them to be $1,5,6$, the carriers read $(0\underline{1}3\underline{56}),(\underline{1}34\underline{56}),(\underline{1}24\underline{56}),(0\underline{1}4\underline{56}),(0\underline{1}3\underline{56})$ from the left to the right, where the underlines signify the added letters. Suppose the added letters are $\alpha,\beta,\gamma$. Then, in the Lagrangian picture mentioned in the above, the effect of the supplement is to endow the RHS of (65) with an extra factor $+\,t^{f_{1}m_{0}+f_{5}(m_{0}+\cdots+m_{4})+f_{6}(m_{0}+\cdots+m_{5})+f_{7}(m_{0}+\cdots+m_{6})}$, where $f_{\lambda}=[\lambda\in\{\alpha,\beta,\gamma\}]=0,1$ and $f_{1}+f_{5}+f_{6}+f_{7}=3$ reflecting that there are three letters to be added. The sign factor is $+$ because a possible -- from any vertex with vertical arrow $h_{q_{i}}\rightarrow h_{q_{i}+1}$ is compensated by the leftmost vertex with vertical arrow $h_{g}\rightarrow 0$. Now, the sum over non-unique carriers for $\dot{T}^{5}(0)$ becomes a sum over the ways to supplement extra letters to the minimal carriers. Consequently we get
@@
$\begin{split}x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{5}(0)|\boldsymbol{\sigma}\rangle_{j}&=x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{2}(0)|\boldsymbol{\sigma}\rangle_{j}\!\sum_{\begin{subarray}{c}f_{1},f_{5},f_{6},f_{7}=0,1\\ f_{1}+f_{5}+f_{6}+f_{7}=3\end{subarray}}t^{f_{1}m_{0}+f_{5}(m_{0}+\cdots+m_{4})+f_{6}(m_{0}+\cdots+m_{5})+f_{7}(m_{0}+\cdots+m_{6})}\\ &=x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{2}(0)|\boldsymbol{\sigma}\rangle_{j}e_{3}(t^{K_{1}},t^{K_{5}},t^{K_{6}},t^{K_{7}}),\end{split}$ (66)
@@
where $K_{i}$ is defined in (7) and $e_{3}$ is an elementary symmetric polynomial (45). In general, a similar argument leads to
$x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}=x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{d}(0)|\boldsymbol{\sigma}\rangle_{j}e_{k-d}(t^{K_{\bar{h}_{1}}},\ldots,t^{K_{\bar{h}_{n-g}}})\qquad(d\leq k\leq n+1),$ (67)
@@
$\{0,\dots,n\}=\{h_{0}(=0),h_{1},\ldots,h_{g}\}\sqcup\{\bar{h}_{1},\ldots,\bar{h}_{n-g}\}.$ (68)
Substituting (65) and (67) into the RHS of (64) and using (8), (68) and (46), we obtain
$\begin{split}D_{\bf m}^{-1}\sum_{k=d}^{n+1}(-1)^{k-1}x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}&=D^{-1}_{\bf m}(1-t)\prod_{i=1}^{g}(1-t)t^{\ell_{h_{i}}}\sum_{k=d}^{n+1}(-1)^{k-d}e_{k-d}(t^{K_{\bar{h}_{1}}},\ldots,t^{K_{\bar{h}_{n-g}}})\\ &=\frac{(1-t)\prod_{i=1}^{g}(1-t)t^{\ell_{h_{i}}}\prod_{i=1}^{n-g}(1-t^{K_{\bar{h}_{i}}})}{(1-t)\prod_{i=1}^{g}(1-t^{K_{h_{i}}})\prod_{i=1}^{n-g}(1-t^{K_{\bar{h}_{i}}})}=\prod_{i=1}^{g}\frac{(1-t)t^{\ell_{h_{i}}}}{1-t^{K_{h_{i}}}},\end{split}$ (69)
completing the proof of (64).
@@ 5.4. Unwanted terms
The unwanted terms correspond to the case where the minimum $h_{0}$ of the moved particle types $(h_{0},\ldots,h_{g})$ in Lemma 5 is nonzero. Thus we are to show
$0=\sum_{k=d}^{n+1}(-1)^{k-1}x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}$ (70)
assuming that the reduced diagram of $x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}$ has the form (59), where $r_{i}$ and $s_{i}$ satisfy the conditions (60a) and (60b) with $h_{0}\in\{1,\ldots,n\}$. All the arguments concerning the wanted terms persist until (67). A key difference arises at (68), where $h_{0}\neq 0$ results in $0\in\{\bar{h}_{1},\ldots,\bar{h}_{n-g}\}$. Since $K_{0}=0$, the summation $\sum_{k=d}^{n+1}(-1)^{k-d}e_{k-d}(t^{K_{\bar{h}_{1}}},\ldots,t^{K_{\bar{h}_{n-g}}})=\prod_{i=1}^{n-g}(1-t^{K_{\bar{h}_{i}}})$ involved in (69) vanishes.
We note that in the above calculation and (69), the summand $e_{k-d}(t^{K_{\bar{h}_{1}}},\ldots,t^{K_{\bar{h}_{n-g}}})$ is actually zero for $k=n+1$, as the index $n+1-d$ exceeds the number $n-g$ of the variables due to $d\leq g$. However, this term is indeed necessary in (55) to ensure that the main formula (51) remains neatly valid, including the diagonal terms. We have thus completed the proof of Theorem 3.
It is natural to consider a generalization of the alternating sum in (51) by introducing a parameter $\zeta$:
$\sum_{k=0}^{n+1}(-\zeta)^{k-1}\dot{T}^{k}(0).$ (71)
Using (65) and (67), we find that its off-diagonal elements take a factorized form:
$\sum_{k=0}^{n+1}(-\zeta)^{k-1}x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}=\zeta^{d-1}(1-t)(1-\zeta t^{K_{\bar{h}_{1}}})\cdots(1-\zeta t^{K_{\bar{h}_{n-g}}})\prod_{i=1}^{g}(1-t)t^{\ell_{h_{i}}},$ (72)
where notation follows (69). In particular for $\zeta=t^{-K_{1}},\ldots,t^{-K_{n}}$, this result reveals an interesting selection rule for nonzero transition coefficients in the process $\boldsymbol{\sigma}\rightarrow\boldsymbol{\sigma}^{\prime}$. However, in general, these coefficients do not satisfy the positivity condition for off-diagonal transition rates.
@@ 6.1. Stationary eigenvalue of $T^{k}(z)$
Let |¯$|\overline{\mathbb{P}}({\bf m})\rangle\in\mathbb{V}({\bf m})$ be the stationary state of the $t$-PushTASEP. It is a unique vector, up to normalization, satisfying $H_{\text{PushTASEP}}|\overline{\mathbb{P}}({\bf m})\rangle=0$. From Theorem 3 and the commutativity (37), it follows that |¯$|\overline{\mathbb{P}}({\bf m})\rangle$ is a joint eigenvector of the transfer matrices $T^{0}(z),\ldots,T^{n+1}(z)$. Moreover, while |¯$|\overline{\mathbb{P}}({\bf m})\rangle$ depends on the inhomogeneities $x_{1},\ldots,x_{L}$, it remains independent of $z$. Let $\Lambda^{k}(z)=\Lambda^{k}(z|x_{1},\ldots,x_{L})$ be the stationary eigenvalue of $T^{k}(z)$, so that $T^{k}(z)|\overline{\mathbb{P}}({\bf m})\rangle=\Lambda^{k}(z)|\overline{\mathbb{P}}({\bf m})\rangle$. Following an analytic Bethe ansatz argument similar to that in [KMMO16, sec. 4.1], we obtain the following expression: [^6] We omit a rigorous derivation in this paper. The result corresponds to the case where all Baxter $Q$ functions become constant, as demonstrated in [KMMO16, Sec. 4.5].
$\Lambda^{k}(z|x_{1},\ldots,x_{L})=e_{k-1}(t^{K_{1}},\ldots,t^{K_{n}})\prod_{j=1}^{L}\left(1-\frac{tz}{x_{j}}\right)+e_{k}(t^{K_{1}},\ldots,t^{K_{n}})\prod_{j=1}^{L}\Bigl{(}1-\frac{z}{x_{j}}\Bigr{)},$ (73)
where $K_{i}$ is defined in (7) and depends on ${\bf m}$. This is a Yang-Baxterization of the $k$’th elementary symmetric polynomial:
$\Lambda^{k}(z)$ $=\prod_{j=1}^{L}d_{k}\Bigl{(}\frac{z}{x_{j}}\Bigr{)}^{-1}\sum_{0\leq i_{1}<\dots<i_{k}\leq n}\framebox{$i_{1}$}_{\,z}\,\framebox{$i_{2}$}_{\,t^{-1}z}\cdots\framebox{$i_{k}$}_{\,t^{-k+1}z},$ (74a)
[…]
where $d_{k}(z)$ is defined by (96). For $k=0$ and $k=n+1$, the formula (73) simplifies to (39) and (40), respectively, as $e_{k-1}(t^{K_{1}},\ldots,t^{K_{n}})$ and $e_{k}(t^{K_{1}},\ldots,t^{K_{n}})$ vanish.
Now let $\dot{\Lambda}(z)=\frac{d\Lambda(z)}{dz}$. Differentiating (73) at $z=0$, we obtain
$\dot{\Lambda}^{k}(0)=-\left(te_{k-1}(t^{K_{1}},\ldots,t^{K_{n}})+e_{k}(t^{K_{1}},\ldots,t^{K_{n}})\right)\sum_{j=1}^{L}\frac{1}{x_{j}}=-e_{k}(t,t^{K_{1}},\ldots,t^{K_{n}})\sum_{j=1}^{L}\frac{1}{x_{j}}.$ (75)
This leads to an interesting interpretation of the quantity $D_{\bf m}$ (8) as
$\sum_{k=0}^{n+1}(-1)^{k-1}\dot{\Lambda}^{k}(0)=\left(\sum_{j=1}^{L}\frac{1}{x_{j}}\right)\sum_{k=0}^{n+1}(-1)^{k}e_{k}(t,t^{K_{1}},\ldots,t^{K_{n}})=\left(\sum_{j=1}^{L}\frac{1}{x_{j}}\right)D_{\bf m}.$ (76)
Consequently, our main formula (51) is also expressed as
$H_{\text{PushTASEP}}(x_{1},\ldots,x_{L})=D_{\bf m}^{-1}\frac{d}{dz}\sum_{k=0}^{n+1}(-1)^{k-1}\left.\left(T^{k}(z|x_{1},\ldots,x_{L})-\Lambda^{k}(z|x_{1},\ldots,x_{L})\right)\right|_{z=0}.$ (77)
@@
$H_{\text{PushTASEP}}(x_{1},\ldots,x_{L})|\overline{\mathbb{P}}({\bf m})\rangle=0$ (78)
@@
$\frac{tx_{1}+x_{3}+tx_{3}}{x_{1}}|012\rangle+\frac{x_{2}+x_{3}+tx_{3}}{x_{2}}|102\rangle+\text{cyc.},$. ,
@@
$\frac{t^{2}x_{1}+x_{4}+tx_{4}+t^{2}x_{4}}{x_{1}}|0112\rangle+\frac{tx_{2}+x_{4}+tx_{4}+t^{2}x_{4}}{x_{2}}|1012\rangle+\frac{x_{3}+x_{4}+tx_{4}+t^{2}x_{4}}{x_{3}}|1102\rangle+\text{cyc.},$. ,
@@
$\begin{split}&\frac{t^{2}x_{1}+t^{2}x_{2}+x_{5}+tx_{5}+t^{2}x_{5}}{x_{1}x_{2}}|00112\rangle+\frac{t^{2}x_{1}+tx_{3}+x_{5}+tx_{5}+t^{2}x_{5}}{x_{1}x_{3}}|01012\rangle\\ &+\frac{tx_{2}+tx_{3}+x_{5}+tx_{5}+t^{2}x_{5}}{x_{2}x_{3}}|10012\rangle+\frac{t^{2}x_{1}+x_{4}+x_{5}+tx_{5}+t^{2}x_{5}}{x_{1}x_{4}}|01102\rangle\\ &+\frac{tx_{2}+x_{4}+x_{5}+tx_{5}+t^{2}x_{5}}{x_{2}x_{4}}|10102\rangle+\frac{x_{3}+x_{4}+x_{5}+tx_{5}+t^{2}x_{5}}{x_{3}x_{4}}|11002\rangle+\text{cyc.}\end{split}$ (79)

### v3

@@ 5.3. Wanted terms
This subsection and the next form the technical focus of the proof. From the definition of $w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)$ around (2.7), the wanted terms generally correspond to the situation where the minimum $h_{0}$ of the moved particle types $(h_{0},\ldots,h_{g})$ in Lemma 5.3 is zero, i.e., $h_{0}=0$. Then (5.15) is written down explicitly as
$\prod_{i=1}^{g}\frac{(1-t)t^{\ell_{h_{i}}}}{1-t^{K_{h_{i}}}}=D_{\bf m}^{-1}\sum_{k=d}^{n+1}(-1)^{k-1}x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}.$ (5.16)
Our calculation of the RHS of (5.16) consists of two steps.
@@
Step 1. We consider the “leading term” $k=d$ in the RHS of (5.15) and the corresponding reduced diagram, in which the carriers are uniquely determined, as shown in Example 5.5 (b) and (c). We claim that
$x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{d}(0)|\boldsymbol{\sigma}\rangle_{j}=(-1)^{d-1}(1-t)\prod_{i=1}^{g}(1-t)t^{\ell_{h_{i}}},$ (5.17)
@@
where $\ell_{h}$ has been defined prior to (2.7). Let us justify the origin of the constituent factors (i) sign, (ii) powers of (1t)(1-t), (iii) powers of $t$, individually.
@@
(i) The sign of a vertex can become negative only for non-diagonal transitions, which occur at the $g+1$ vertices in the reduced diagram (5.9). From the comment following (3.12) and the conditions in Lemma 5.3, the $g+1$ vertices corresponding to the vertical arrows $r_{i}\rightarrow s_{i}$ in (5.10b) have ++ signs for $i=1,\ldots,g$ and $(-1)^{d-1}$ for $i=0$.
(ii) From Table 1, the contributions of (1t)(1-t) at each of the $g+1$ vertices results in a factor of $(1-t)^{g+1}$.
@@
(iii) From Table 1, the power of $t$ can be evaluated as the sum of the quantities of the form $a_{c+1}+\cdots+a_{n}$ in the multiplicity representation of the carriers ${\bf a}=(a_{0},\ldots,a_{n})$, attached to each vertex. This formula implies that a particle of type $h$ in the carriers contributes $c<h]\in\{0,1\}$ whenever it passes over a site $i$ occupied with $\sigma_{i}=c$. Alternatively, this can be calculated as the total contribution collected by the moved particles $h_{0},\ldots,h_{g}$ from the smaller-species particles in $\boldsymbol{\sigma}$. This precisely leads to $\ell_{h_{1}}+\cdots+\ell_{h_{g}}$, where $\ell_{h_{0}}$ can be excluded due to $\ell_{h_{0}}=\ell_{0}=0$. Thus the factor $\prod_{i=1}^{g}t^{\ell_{h_{i}}}$ is obtained as claimed. The reformulation in the calculation described here is analogous to the transition from the Eulerian picture, which tracks properties at fixed spatial points, to the Lagrangian picture, which follows individual particles, in fluid mechanics. In our context, it also incorporates the contribution from the vertices corresponding to the diagonal transitions efficiently via the quantities $\ell_{h}$’s.
@@
Step 2. Let us turn to the $k=d+1,\ldots,n+1$ terms in (5.15). We illustrate the idea of evaluating them along Example 5.5 (b) for $k=5$ and $n=7$ $(d=2,g=3)$. The carriers from $\mathscr{B}^{5}$ are no longer unique. However, those satisfying the weight conservation with $\boldsymbol{\sigma}$ and $\boldsymbol{\sigma}^{\prime}$ are exactly those obtained just by supplementing the common three letters from the yet unused ones $\{1,5,6,7\}$ to the existing ones everywhere. For instance, choosing them to be $1,5,6$, the carriers read $(0\underline{1}3\underline{56}),(\underline{1}34\underline{56}),(\underline{1}24\underline{56}),(0\underline{1}4\underline{56}),(0\underline{1}3\underline{56})$ from the left to the right, where the underlines signify the added letters. Suppose the added letters are $\alpha,\beta,\gamma$. Then, in the Lagrangian picture mentioned in the above, the effect of the supplement is to endow the RHS of (5.17) with an extra factor $t^{f_{1}m_{0}+f_{5}(m_{0}+\cdots+m_{4})+f_{6}(m_{0}+\cdots+m_{5})+f_{7}(m_{0}+\cdots+m_{6})}$, where $f_{\lambda}=[\lambda\in\{\alpha,\beta,\gamma\}]=0,1$ and $f_{1}+f_{5}+f_{6}+f_{7}=3$ reflecting that there are three letters to be added. The sign factor is ++ because a possible - from any vertex with vertical arrow $h_{q_{i}}\rightarrow h_{q_{i}+1}$ is compensated by the leftmost vertex with vertical arrow $h_{g}\rightarrow 0$. Now, the sum over non-unique carriers for $\dot{T}^{5}(0)$ becomes a sum over the ways to supplement extra letters to the minimal carriers. Consequently we get
@@
$\begin{split}x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{5}(0)|\boldsymbol{\sigma}\rangle_{j}=&x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{2}(0)|\boldsymbol{\sigma}\rangle_{j}\\ &\times\!\sum_{\begin{subarray}{c}f_{1},f_{5},f_{6},f_{7}=0,1\\ f_{1}+f_{5}+f_{6}+f_{7}=3\end{subarray}}t^{f_{1}m_{0}+f_{5}(m_{0}+\cdots+m_{4})+f_{6}(m_{0}+\cdots+m_{5})+f_{7}(m_{0}+\cdots+m_{6})}\\ =&x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{2}(0)|\boldsymbol{\sigma}\rangle_{j}e_{3}(t^{K_{1}},t^{K_{5}},t^{K_{6}},t^{K_{7}}),\end{split}$ (5.18)
@@
where $K_{i}$ is defined in (2.3) and $e_{3}$ is an elementary symmetric polynomial (4.11). In general, a similar argument leads to
$x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}=x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{d}(0)|\boldsymbol{\sigma}\rangle_{j}e_{k-d}(t^{K_{\bar{h}_{1}}},\ldots,t^{K_{\bar{h}_{n-g}}})\qquad(d\leq k\leq n+1),$ (5.19)
@@
$\{0,\dots,n\}=\{h_{0}(=0),h_{1},\ldots,h_{g}\}\sqcup\{\bar{h}_{1},\ldots,\bar{h}_{n-g}\}.$ (5.20)
Substituting (5.17) and (5.19) into the RHS of (5.16) and using (2.4), (5.20) and (4.12), we obtain
$\begin{split}\sum_{k=d}^{n+1}\frac{(-1)^{k-1}}{D_{\bf m}}x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}&=\frac{(1-t)}{D_{\bf m}}\prod_{i=1}^{g}(1-t)t^{\ell_{h_{i}}}\sum_{k=d}^{n+1}(-1)^{k-d}e_{k-d}(t^{K_{\bar{h}_{1}}},\ldots,t^{K_{\bar{h}_{n-g}}})\\ &=\frac{(1-t)\prod_{i=1}^{g}(1-t)t^{\ell_{h_{i}}}\prod_{i=1}^{n-g}(1-t^{K_{\bar{h}_{i}}})}{(1-t)\prod_{i=1}^{g}(1-t^{K_{h_{i}}})\prod_{i=1}^{n-g}(1-t^{K_{\bar{h}_{i}}})}=\prod_{i=1}^{g}\frac{(1-t)t^{\ell_{h_{i}}}}{1-t^{K_{h_{i}}}},\end{split}$ (5.21)
completing the proof of (5.16).
@@ 5.4. Unwanted terms
The unwanted terms correspond to the case where the minimum $h_{0}$ of the moved particle types $(h_{0},\ldots,h_{g})$ in Lemma 5.3 is nonzero. Thus we are to show
$0=\sum_{k=d}^{n+1}(-1)^{k-1}x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}$ (5.22)
assuming that the reduced diagram of $x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}$ has the form (5.9), where $r_{i}$ and $s_{i}$ satisfy the conditions (5.10a) and (5.10b) with $h_{0}\in\{1,\ldots,n\}$. All the arguments concerning the wanted terms persist until (5.19). A key difference arises at (5.20), where $h_{0}\neq 0$ results in $0\in\{\bar{h}_{1},\ldots,\bar{h}_{n-g}\}$. Since $K_{0}=0$, the summation $\sum_{k=d}^{n+1}(-1)^{k-d}e_{k-d}(t^{K_{\bar{h}_{1}}},\ldots,t^{K_{\bar{h}_{n-g}}})=\prod_{i=1}^{n-g}(1-t^{K_{\bar{h}_{i}}})$ involved in (5.21) vanishes.
We note that in the above calculation and (5.21), the summand $e_{k-d}(t^{K_{\bar{h}_{1}}},\ldots,t^{K_{\bar{h}_{n-g}}})$ is actually zero for $k=n+1$, as the index n+1dn+1-d exceeds the number ngn-g of the variables due to $d\leq g$. However, this term is indeed necessary in (5.5) to ensure that the main formula (5.1) remains neatly valid, including the diagonal terms. We have thus completed the proof of Theorem 5.1.
It is natural to consider a generalization of the alternating sum in (5.1) by introducing a parameter $\zeta$:
$\sum_{k=0}^{n+1}(-\zeta)^{k-1}\dot{T}^{k}(0).$ (5.23)
Using (5.17) and (5.19), we find that its off-diagonal elements take a factorized form:
$\sum_{k=0}^{n+1}(-\zeta)^{k-1}x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}=\zeta^{d-1}(1-t)(1-\zeta t^{K_{\bar{h}_{1}}})\cdots(1-\zeta t^{K_{\bar{h}_{n-g}}})\prod_{i=1}^{g}(1-t)t^{\ell_{h_{i}}},$ (5.24)
where notation follows (5.21). In particular for $\zeta=t^{-K_{1}},\ldots,t^{-K_{n}}$, this result reveals an interesting selection rule for nonzero transition coefficients in the process $\boldsymbol{\sigma}\rightarrow\boldsymbol{\sigma}^{\prime}$. However, in general, these coefficients do not satisfy the positivity condition for off-diagonal transition rates.
@@ 6.1. Stationary eigenvalue of $T^{k}(z)$
Let $|\overline{\mathbb{P}}({\bf m})\rangle\in\mathbb{V}({\bf m})$ be the stationary state of the $t$-PushTASEP. It is a unique vector, up to normalization, satisfying $H_{\text{PushTASEP}}|\overline{\mathbb{P}}({\bf m})\rangle=0$. From Theorem 5.1 and the commutativity (4.3), it follows that $|\overline{\mathbb{P}}({\bf m})\rangle$ is a joint eigenvector of the transfer matrices $T^{0}(z),\ldots,T^{n+1}(z)$. Moreover, while $|\overline{\mathbb{P}}({\bf m})\rangle$ depends on the inhomogeneities $x_{1},\ldots,x_{L}$, it remains independent of $z$. Let $\Lambda^{k}(z)=\Lambda^{k}(z|x_{1},\ldots,x_{L})$ be the stationary eigenvalue of $T^{k}(z)$, so that $T^{k}(z)|\overline{\mathbb{P}}({\bf m})\rangle=\Lambda^{k}(z)|\overline{\mathbb{P}}({\bf m})\rangle$. Following an analytic Bethe ansatz argument similar to that in [KMMO16, sec. 4.1], we obtain the following expression: [^5] We omit a rigorous derivation in this paper. The result corresponds to the case where all Baxter $Q$ functions become constant, as demonstrated in [KMMO16, Sec. 4.5].
$\Lambda^{k}(z|x_{1},\ldots,x_{L})=e_{k-1}(t^{K_{1}},\ldots,t^{K_{n}})\prod_{j=1}^{L}\left(1-\frac{tz}{x_{j}}\right)+e_{k}(t^{K_{1}},\ldots,t^{K_{n}})\prod_{j=1}^{L}\Bigl(1-\frac{z}{x_{j}}\Bigr),$ (6.1)
where $K_{i}$ is defined in (2.3) and depends on $\bf m$}. This is a Yang-Baxterization of the $k$’th elementary symmetric polynomial:
$\Lambda^{k}(z)$ $=\prod_{j=1}^{L}d_{k}\Bigl(\frac{z}{x_{j}}\Bigr)^{-1}\sum_{0\leq i_{1}<\dots<i_{k}\leq n}\framebox{$i_{1}$}_{\,z}\,\framebox{$i_{2}$}_{\,t^{-1}z}\cdots\framebox{$i_{k}$}_{\,t^{-k+1}z},$ (6.2a)
$\framebox{$i$}_{\,z}$ $=t^{K_{i}}\prod_{j=1}^{L}\left(1-t^{\delta_{i,0}}\frac{z}{x_{j}}\right),$ (6.2b)
where $d_{k}(z)$ is defined by (3.8). For $k=0$ and $k=n+1$, the formula (6.1) simplifies to (4.5) and (4.6), respectively, as $e_{k-1}(t^{K_{1}},\ldots,t^{K_{n}})$ and $e_{k}(t^{K_{1}},\ldots,t^{K_{n}})$ vanish.
Now let $\dot{\Lambda}(z)=\frac{d\Lambda(z)}{dz}$. Differentiating (6.1) at $z=0$, we obtain
$\dot{\Lambda}^{k}(0)=-\left(te_{k-1}(t^{K_{1}},\ldots,t^{K_{n}})+e_{k}(t^{K_{1}},\ldots,t^{K_{n}})\right)\sum_{j=1}^{L}\frac{1}{x_{j}}=-e_{k}(t,t^{K_{1}},\ldots,t^{K_{n}})\sum_{j=1}^{L}\frac{1}{x_{j}}.$ (6.3)
This explains the origin of the quantity $D_{\bf m}$ in (2.4) in the derivative of eigenvalues as
$\sum_{k=0}^{n+1}(-1)^{k-1}\dot{\Lambda}^{k}(0)=\left(\sum_{j=1}^{L}\frac{1}{x_{j}}\right)\sum_{k=0}^{n+1}(-1)^{k}e_{k}(t,t^{K_{1}},\ldots,t^{K_{n}})=\left(\sum_{j=1}^{L}\frac{1}{x_{j}}\right)D_{\bf m}.$ (6.4)
Consequently, our main formula (5.1) is also expressed as
$H_{\text{PushTASEP}}(x_{1},\ldots,x_{L})=D_{\bf m}^{-1}\frac{d}{dz}\sum_{k=0}^{n+1}(-1)^{k-1}\left.\left(T^{k}(z|x_{1},\ldots,x_{L})-\Lambda^{k}(z|x_{1},\ldots,x_{L})\right)\right|_{z=0}.$ (6.5)
@@
$H_{\text{PushTASEP}}(x_{1},\ldots,x_{L})|\overline{\mathbb{P}}({\bf m})\rangle=0$ (6.6)
@@
$\frac{tx_{1}+x_{3}+tx_{3}}{x_{1}}|012\rangle+\frac{x_{2}+x_{3}+tx_{3}}{x_{2}}|102\rangle+\text{cyc.},$
@@
$\frac{t^{2}x_{1}+x_{4}+tx_{4}+t^{2}x_{4}}{x_{1}}|0112\rangle+\frac{tx_{2}+x_{4}+tx_{4}+t^{2}x_{4}}{x_{2}}|1012\rangle\\ +\frac{x_{3}+x_{4}+tx_{4}+t^{2}x_{4}}{x_{3}}|1102\rangle+\text{cyc.},$
@@
$\frac{t^{2}x_{1}+t^{2}x_{2}+x_{5}+tx_{5}+t^{2}x_{5}}{x_{1}x_{2}}|00112\rangle+\frac{t^{2}x_{1}+tx_{3}+x_{5}+tx_{5}+t^{2}x_{5}}{x_{1}x_{3}}|01012\rangle\\ +\frac{tx_{2}+tx_{3}+x_{5}+tx_{5}+t^{2}x_{5}}{x_{2}x_{3}}|10012\rangle+\frac{t^{2}x_{1}+x_{4}+x_{5}+tx_{5}+t^{2}x_{5}}{x_{1}x_{4}}|01102\rangle\\ +\frac{tx_{2}+x_{4}+x_{5}+tx_{5}+t^{2}x_{5}}{x_{2}x_{4}}|10102\rangle+\frac{x_{3}+x_{4}+x_{5}+tx_{5}+t^{2}x_{5}}{x_{3}x_{4}}|11002\rangle+\text{cyc.}$